use method of undetermined coefficients to solve ivp

y" - 4y' - 12y = 3e^5x, y(0) = 18/7, y'(0) = -1/7
Answer #1

y-44-lay = 3e55 ) yo) = yko) = - $ Complementory Solution , f_4y_124 (D9-4D-12)4=0 CO9-40-12)=0 (D+2) CD-6)=0 D -- 2,6 foq=&typ - a eextca ega $ 5x yo) 18 T =crtca - 34 Gtca = 3 O y! -20, ēdze 6x 5520 y co) = + +6C2 = -26,76ca - 1Ş >> -26+6C2 = 2